# Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs

### v2

@@ Abstract
Linguistic evaluations of how well LMs generalize to produce or understand novel text often implicitly take for granted that natural languages are generated by symbolic rules. Grammaticality is thought to be determined by whether sentences obey such rules. Interpretation is believed to be compositionally generated by syntactic rules operating on meaningful words. Semantic parsing is intended to map sentences into formal logic. Failures of LMs to obey strict rules have been taken to reveal that LMs do not produce or understand language like humans. Here we suggest that LMs’ failures to obey symbolic rules may be a feature rather than a bug, because natural languages are not based on rules. New utterances are produced and understood by a combination of flexible, interrelated, and context-dependent constructions. We encourage researchers to reimagine appropriate benchmarks and analyses that acknowledge the rich, flexible generalizations that comprise natural languages.
@@ 1 Introduction
Constructions: Learned Pairings of Form and Function Rules
(Partially-filled) words, common and rare (partially-filled) schemata Common abstract patterns
Wide range of functions Only abstract functions
Combination of open slots and/or fixed lexical units Include only variables
Context-sensitive (and plentiful) Context-free (and few)
Inter-related within a complex network Unstructured list
Sensitive to similarity and frequency Insensitive to similarity or frequency
Slots constrained in open-ended range of ways Variables constrained by gram. category
Table 1: Differences between constructions and rules
@@
How well do large Language Models (LMs) generalize beyond their training data? Much work on this question has presumed that generalizations require symbolic rules for syntax and semantics that generate acceptable new forms and compositional meanings. Rules are invoked to explain that if you learn a new modifier (‘blonky’) and a new count noun (‘gravimin’), a compositional rule could predict that ‘a blonky gravimin’ is a gravimin that is blonky. In what follows, we use “rule” to refer to context-free generalizations that contain variables, to be instantiated by any instance of a general type. In the domain of language, rules are assumed to be uninfluenced by frequency, similarity, or context Pinker (1999). Rules of this type are relied on in math, logic, and programming languages.
@@
Because earlier statistical models (e.g., n-gram or Markov models) seemed unable to generalize fully or capture non-local dependencies Chomsky (1957), rules seemed to many to be the only game in town for human language. After all, if a standard bigram model hadn’t seen ‘blonky gravimin’ before, it would be unable to interpret it. Influential thinkers argued that neural networks, which did not involve rules, would never be appropriate models of human cognition for this reason Fodor and Pylyshyn (1988); Pinker and Prince (1988); Marcus (1998); Fodor and Lepore (2002); Marcus (2001); Calvo and Symons (2014). Yet GPT-4o, for example, not only recognizes ‘a blonky gravimin’ is a noun phrase, it explicitly guesses several naturalistic interpretations, e.g., ‘A person or act that awkwardly and absurdly pretends to be serious.’
@@
Critically, today’s LMs arose from statistical, distributional parallel models Mikolov et al. (2013); Rumelhart et al. (1986) rather than rule-based natural language technologies. They do not rely on hard-coded rules, yet their ability to produce coherent, naturalistic language and respond appropriately is unparalleled by purely symbolic systems Piantadosi (2024); Goldberg (2024); Weissweiler et al. (2023); Hofmann et al. (2025).
@@
Nonetheless, an assumption that generalizations are equivalent to rules continues to motivate many evaluations of syntax, meaning, and their compositional combination: e.g., Natural Language Inference Bowman et al. (2015), Semantic Parsing Palmer et al. (2005); Reddy et al. (2017), tests of binary grammatical acceptability Warstadt et al. (2019); Dentella et al. (2023) and rule-based compositionality Kim and Linzen (2020). Together, such tasks made up more than half of the GLUE benchmark Wang et al. (2018), created to evaluate language models on their skill at being “general, flexible, and robust.” Lackluster performance on rule-based tasks in the early days of LMs was taken to imply that the models did not use language the way people do and were instead merely imitating shallow surface patterns Lake and Baroni (2018); Kim and Linzen (2020); Weißenhorn et al. (2022); Bolhuis et al. (2023). In a survey of 79 NLP researchers, McCurdy et al. (2024) reported that 87% believed LMs were not sufficiently compositional and a sizable proportion (39%) believed explicit discrete symbolic rules were required.
@@
Evaluations of early LMs’ challenges using algebraic or logical rules did expose certain shortcomings in their ability to reason abstractly and solve math problems (e.g., Mahowald et al., 2024). At the same time, LM’s ability to produce and respond to natural languages naturalistically is hard to overstate (e.g., Coil and Shwartz, 2023). We therefore suggest that rule-based evaluations have been over-emphasized in the domain of natural language production and comprehension. More nuanced metrics, measuring gradient judgments and context-dependent interpretations are more valuable (e.g., Juzek, 2024; Hu et al., 2024a).
Rules are not sufficient for generalization. For LMs to use language like humans, they require interpretations that are far richer than rules can provide for thousands of collocations, conventional metaphors, idioms, and context-dependent interpretations. Abstract grammatical patterns likewise regularly convey semantic and/or pragmatic information that restrict their contexts of use and interpretation.
@@
Just as LMs appear to do Tseng et al. (2022), we propose that each natural language requires mastering a network of hundreds of thousands of context-dependent, gradient, flexible schemata (constructions, see §5), which often contain ‘slots’ that constrain their fillers and how those fillers are interpreted. Constrained slots allow for new combinations, flexibly adapted in context. For instance, the phrase ’<time period> ago’ can coerce a temporal interpretation of filler phrases that do not designate time periods (e.g., ‘three rest stops ago’). Rather than rule-based composition of meaning, composition-by-construction allows constructions to contribute meaningfully to interpretation in ways that range from quite specific to largely abstract.
A list of differences between rules and constructions is provided in Table 1, and we explain each row in §5. Constructions anticipate that different languages or dialects will provide speakers with different ConstructionNets, so that cross-linguistic differences can be captured naturally.
@@
We feel that context-free, fully abstract rules are not necessary, as we suggest in §5 that constructions do what rules are meant to do and more. At the same time, the approach we advocate need not eschew the possibility of context-free, fully abstract rules that apply within a particular language, or even, in the limiting case, across languages. If such cases exist, they can be treated as a fully abstract construction Jackendoff (2002). Note that Jackendoff (2002) proposes a rule Verb + Particle to capture the syntax of English complex verbs, but no one would argue that the meanings of complex verbs arise from a compositional rule (e.g., one can look up a number or look down on someone but not ?look up on someone or ?look down a number). We are simply advocating for an increased focus on the extent to which and how LMs manage to produce and comprehend human-like natural languages in all their context-specificity and complexity.
@@
Many of our theoretical points are not new. While early AI relied on algebraic rules Minsky and Papert (1969); Lenat (1995), many researchers soon realized that rules were too brittle to scale up beyond highly restricted domains such as artificial block worlds Winograd (1980). Neural network researchers have continuously argued in favor of a single representational system and against the usefulness of rules, primarily in the domain of words and inflectional morphology (e.g., Rumelhart et al., 1986; Rogers and McClelland, 2004; Elman, 2009; Christiansen and Chater, 1999; MacDonald et al., 1994; McClelland, 2015).
Our contribution is to review leading paradigms used in LM evaluations of syntax (§2), semantics (§3), and compositionality (§4). We explain how rules are implicitly assumed in many impactful evaluations, briefly describing how the assumptions arose, and why they are problematic. We propose an alternative theoretical basis (§5), encouraging the field to evaluate LMs on the extent to which they learn and represent the complex network of constructions that comprises each language and how constructions are combined and generalized to produce new messages in new contexts.
@@ 2 Syntactic Rules in LM Evaluations
An assumption that natural languages are generated by rules such as phrase structure rules, movement rules, or an operation ‘Merge’ has been assumed by most versions of generative grammar since Chomsky (1957). Syntactic rules are intended to operate on broad and clearly defined grammatical categories (e.g., nouns, verbs, adjectives) or relations (e.g., subject, object), which are in turn combined in rule-like fashion to create larger units (e.g., noun phrases, verb phrases, adjective phrases). Rules are claimed to be uninfluenced by frequency, similarity, meaning, or context Pinker (1999).
@@
Human judgments on sentences are gradient rather than binary, and demonstrably depend on frequency, plausibility, complexity, memory demands, potential alternatives, and context Grodner and Gibson (2005); Schütze and Sprouse (2013); Robenalt and Goldberg (2015); Gibson and Hickok (1993); Fang et al. (2023). The amount of exposure to written language or linguistic theory also influences judgments. For instance, Dąbrowska (2010) found that laypeople’s judgments on sentences containing long-distance dependencies were more sensitive to lexical content than linguists’ judgments were.
Nonetheless, LMs’ language skills are often evaluated on binary grammaticality judgments on sentences Dentella et al. (2024, 2023); Warstadt et al. (2019) Even sentences included in linguistic textbooks, which one might presume to have clear-cut judgments, in reality are judged gradiently by people Juzek (2024).
The fact that human judgments are gradient can have profound consequences on evaluations. For instance, Dentella et al. (2023) compared humans and LMs against predetermined binary acceptability labels, reporting that LMs’ performance correlated poorly. However, comparing gradient perplexity measures with the same human judgments revealed a strong positive correlation Hu et al. (2024b). Using perplexity measures for models (as well as allowing humans to provide ordinal or gradient judgments) is a step in the right direction Hu et al. (2024b); Juzek (2024).
@@
Parsing text for universal dependencies (UD, de Marneffe et al., 2021) has become a well-established task for evaluating models Zeman et al. (2017, 2018), and since Hewitt and Manning (2019) showed BERT Devlin et al. (2019) to be somewhat skilled in UD, it has became the default operationalization of syntax in the NLP world Amini et al. (2023); Kryvosheieva and Levy (2025); Müller-Eberstein et al. (2022) and in discussions of inductive biases Lindemann et al. (2024); Glavaš and Vulić (2021). UD annotations are partially determined by semantics, and they are based on lexical items, which makes them closer to the approach advocated here rather than abstract phrase structure rules. However, UD analyses presume a universal set of grammatical relations, which is problematic, because not all languages employ the same constructs. That is, there is no universally valid way to define or test for nouns, verbs, subjects, or direct objects (e.g., Croft, 2001). Moreover, UD requires an asymmetric relationship between a ’head’ and dependent, yet the long tail of language includes headless constructions (e.g., the Xer, the Yer construction) Michaelis (2003) and co-headed constructions (e.g., phrasal verbs, conjunctions, and idioms). Therefore, evaluating LMs on UD annotations may set a standard that is misaligned wth aspects of natural languages. Most importantly, LMs produce and interpret meaningful language to varying degrees. Yet, since UD is a syntactic annotation schema, it provides no means of generating meaningful language nor assigning meaningful interpretations.
@@ 3 Semantic Rules in LM Evaluation
Formal logic was developed as a branch of mathematics, used to prove mathematical and philosophical theorems and identify provability gaps Frege (1918); Russell (1905); Gödel (1931). It was based on algebraic rules operating on clear-cut and broadly defined categories. Notably, logicians did not generally assume nor endorse using formal logic to represent the meanings of natural language utterances Carnap (1937); Baker and Hacker (1986), recognizing that natural languages differ from formal logic in many ways. For instance, logic treats and and but as equivalent. It does not provide a natural way to capture commands or questions (Austin, 1975), nor does it naturally distinguish presuppositions from assertions Strawson (1967). Finally, formal logic is not intended to capture effects of context Wittgenstein (1953); Russin et al. (2024).
@@
Natural Language Inference tasks label the second of two sentences as an entailment, contradiction, or neutral. NLI tasks were originally used to train models Wang et al. (2019); Dagan et al. (2006); Nie et al. (2020). Today, they are used as zero-shot evaluation metrics to assess natural language understanding Zhou et al. (2024); McCoy et al. (2019). Entailment and contradiction are key concepts in mathematical proofs. However, the centrality of entailment and contradiction in language understanding can be overstated (e.g., “The semantic concepts of entailment and contradiction are central to all aspects of natural language meaning. ” Bowman et al. (2015) ) (emphasis added, see also Katz, 1972; van Benthem, 2008). Necessary and plausible inferences are a critical aspect of natural language understanding, but they are highly dependent on context, as well as the interlocutors’ general communicative goals.
Humans’ goal is to make sense of others’ messages, so we assume others are trying to be relevant and helpful and do our best to assign coherent meanings to all utterances Grice (1975). For example, outside of logic classes or heated arguments, people rarely conclude that two statements made by the same person are contradictory. If someone utters: ‘The boy is depressed but he is not DEPRESSED’, listeners do not throw up their hands and shout ‘contradiction!’. Instead, they may infer that the boy in question is only somewhat, and not extremely, depressed, or ask to learn more. Humans also assign distinct context-dependent interpretations to apparent tautologies such as ‘Either it’s alive or it’s not’ and ‘If it snows, it snows.’ NLI tasks that rely on judging contradictions or entailments may over- or under-estimate how well LMs’ understanding of natural language aligns with humans’, particularly when binary judgments are required (cf. Dentella et al., 2024).
We need evaluation metrics to take context and communicative goals into account and to allow to allow gradient judgments. An example of the type of evaluation we endorse can be found in the underappreciated NOPE testbed created by Parrish et al. (2021). The authors selected 10 distinct constructions that trigger presuppositions, and curated 100+ instances of each one, based on naturally occurring examples. Each stimulus included two preceding sentences for context. The authors then collected gradient judgments from human raters that allowed them to use their ‘background knowledge about how the world works’ and compared the accuracy of several models, with several controls in place, which strikes us as the right approach.
@@
Banarescu et al. (2013) introduced abstract graphical meaning representations (AMR) for sentential meaning that importantly includes aspects of lexical semantics. It was created to offer a repository of structured meanings to be used for evaluating understanding in LMs, a worthwhile goal (Li et al., 2023; Qiu et al., 2022; Shaw et al., 2021, see also §4). At the same time, the representation explicitly ignores ’syntactic idiosyncracies.’ For example, ‘he described her as a genius’ and ‘his description of her: genius’ are assigned the same AMR. Yet the former is unambiguously about her intellect, while the latter may instead be used to compliment his cleverly tactful description. More generally, simplifications of distinctions made in a natural language commonly result in a loss in meaning, since two utterances are rarely fully interchangeable. Languages rarely offer two ways to express the identical message in context.

### v3

@@ Abstract
Linguistic evaluations of how well LMs generalize to produce or understand language often implicitly take for granted that natural languages are generated by symbolic rules. According to this perspective, grammaticality is determined by whether sentences obey such rules. Interpretation is compositionally generated by syntactic rules operating on meaningful words. Semantic parsing maps sentences into formal logic. Failures of LMs to obey strict rules are presumed to reveal that LMs do not produce or understand language like humans. Here we suggest that LMs’ failures to obey symbolic rules may be a feature rather than a bug, because natural languages are not based on neatly separable, compositional rules. Rather, new utterances are produced and understood by a combination of flexible, interrelated, and context-dependent constructions. Considering gradient factors such as frequencies, context, and function will help us reimagine new benchmarks and analyses to probe whether and how LMs capture the rich, flexible generalizations that comprise natural languages.
@@ 1 Introduction
How well do large Language Models (LMs) generalize beyond their training data? Much work on this question has presumed that generalizations require symbolic rules for syntax and semantics that generate acceptable new forms and compositional meanings. Rules are invoked to explain that if you learn a new modifier (‘blonky’) and a new count noun (‘gravimin’), a compositional rule could predict that ‘a blonky gravimin’ is a gravimin that is blonky. In what follows, we use “rule” to refer to context-free generalizations that contain variables, to be instantiated by any instance of a general type, uninfluenced by frequency, similarity, or context Pinker (1999). Our focus here is on the use of a strict algebraic conception of rules, which we argue, underlies certain approaches to NLP evaluation, even though the notion of a rule is used variably in linguistics today, with several frameworks incorporating functional and/or frequency-based attributes into representations (e.g., Bresnan et al., 2007; Brehm et al., 2022; O’Donnell, 2015),
@@
Because early statistical models (e.g., n-gram or Markov models) seemed unable to generalize fully or capture non-local dependencies Chomsky (1957), early on, rules seemed to many to be the only game in town for human language. After all, if a standard bigram model hadn’t seen ‘blonky gravimin’ before, it would be unable to form a representation of it. Influential thinkers argued that neural networks, which did not involve rules, would never be appropriate models of human cognition for this reason Fodor and Pylyshyn (1988); Pinker and Prince (1988); Marcus (1998); Fodor and Lepore (2002); Marcus (2001); Calvo and Symons (2014).
@@
However, current LMs arose from statistical, distributional parallel models Mikolov et al. (2013); Rumelhart et al. (1986) rather than rule-based natural language technologies. They do not rely on hard-coded rules, yet their ability to produce coherent, naturalistic language and respond appropriately is unparalleled by purely symbolic systems Piantadosi (2024); Goldberg (2024); Weissweiler et al. (2023); Hofmann et al. (2025). GPT-4o, for example, not only recognizes ‘a blonky gravimin’ as a noun phrase, it explicitly offers several naturalistic interpretations, e.g., ‘A person or act that awkwardly and absurdly pretends to be serious.’
@@
Nonetheless, an assumption that generalizations are equivalent to rules continues to motivate many evaluations of syntax, meaning, and their compositional combination: e.g., Natural Language Inference Bowman et al. (2015), Semantic Parsing Palmer et al. (2005); Reddy et al. (2017), tests of binary grammatical acceptability Warstadt et al. (2019); Dentella et al. (2023) and rule-based compositionality Kim and Linzen (2020). Together, such tasks made up more than half of the GLUE benchmark Wang et al. (2018), created to evaluate language models on their skill at being “general, flexible, and robust.” Lackluster performance on rule-based tasks in the early days of LMs was taken to imply that the models did not use language the way people do and were instead merely imitating shallow surface patterns Bender and Koller (2020); Kim and Linzen (2020); Weißenhorn et al. (2022); Bolhuis et al. (2023). In a survey of 79 NLP researchers, McCurdy et al. (2024) reported that 87% believed LMs were not sufficiently compositional and a sizable proportion (39%) believed explicit discrete symbolic rules were required.
@@
Evaluations of LMs’ early challenges with algebraic or logical rules did expose certain shortcomings in their ability to reason abstractly and solve math problems (see e.g., Mahowald et al., 2024). At the same time, LM’s concurrent ability to produce and respond to natural languages naturalistically is hard to overstate (e.g., Coil and Shwartz, 2023).
@@
Mastering a natural language requires mastering a network of hundreds of thousands of context-dependent, gradient, flexible schemata (constructions, see §5), which often contain ‘slots’ that constrain their fillers and how those fillers are interpreted. Constrained slots allow for new combinations, flexibly adapted in context. For instance, the phrase ’<time period> ago’ can coerce a temporal interpretation of filler phrases that do not designate time periods (e.g., ‘three rest stops ago’). Rather than rule-based compositionality, composition-by-construction allows constructions to contribute meaningfully to interpretation in ways that range from abstract to quite narrow and specific. Therefore, for LMs to use language like humans, they require interpretations that are far richer than rules can provide for thousands of collocations, conventional metaphors, idioms, and context-dependent interpretations. Even abstract grammatical patterns also regularly convey semantic and/or pragmatic information that restrict their contexts of use and interpretations. Since different languages and dialects provide speakers with different networks of constructions (ConstructionNets), cross-linguistic differences can be captured naturally.
We suggest that rule-based evaluations have been over-emphasized in the domain of natural language production and comprehension. Our goal is to emphasize the importance of recognizing context, frequencies, meaning and other gradient functional factors in modern evaluations of natural language.
@@
We do not argue that no categorical rule exists in any language. If a categorical rule is needed, it can be treated as the limiting case, a fully abstract construction Jackendoff (2002). For instance, Jackendoff (2002) proposes a symbolic Verb + Particle rule for the syntax of English complex verbs. At the same time, the meanings of individual verb plus particle combinations are far from compositional by any general rule (e.g., one can look up a number or look down on someone but not ?look up on someone nor ?look down a number). Here we advocate for an increased focus on the extent to which and how LMs manage to produce and comprehend human-like natural languages in all their context-specificity and complexity.
@@
Many of our theoretical points are not new, particularly in the domain of morphology. Neural network researchers have continuously argued in favor of a single representational system and against the usefulness of rules in the domain of words and inflectional morphology (e.g., Rumelhart et al., 1986; Rogers and McClelland, 2004; Elman, 2009; Christiansen and Chater, 1999; MacDonald et al., 1994; McClelland, 2015). While early work in Artificial Intelligence relied on algebraic rules Minsky and Papert (1969); Lenat (1995), many researchers soon realized that rules were too brittle to scale up beyond highly restricted domains such as artificial block worlds Winograd (1980).
Our contribution is to review leading paradigms used in LM evaluations of syntax (§2), semantics (§3), and compositionality (§4). We argue that, while these paradigms have been fruitful, they inherit from a tradition that was overly focused on rules, hierarchy, compositionality, and a binary notion of grammaticality. We briefly characterize how these assumptions arose and how they are baked into evaluations. We argue that evaluations that presume categorical and strictly compositional language ignore some of the richest elements of human language. We review construction-based and gradient functionalist approaches to language, arguing that this tradition points to certain lacunae in existing evaluations and open up new possibilities for evaluating natural language understanding. Early work in this direction has already included more nuanced metrics, measuring gradient judgments and context-dependent interpretations (e.g., Juzek, 2024; Hu et al., 2024).
@@ 2 Syntactic Rules in LM Evaluations
Evaluations of the syntactic capabilities of LMs have frequently assumed a binary categorical notion of grammaticality, which is then used to create datasets for evaluation. Below, we discuss several such cases, attempting to make these assumptions explicit to show their limitations.
@@
Human judgments on sentences are gradient rather than binary, and demonstrably depend on frequency, plausibility, complexity, memory demands, potential alternatives, and context Grodner and Gibson (2005); Schütze and Sprouse (2013); Robenalt and Goldberg (2015); Gibson and Hickok (1993); Fang et al. (2023). The amount of exposure to written language or linguistic theory also influences people’s judgments. For instance, Dąbrowska (2010) found that laypeople’s judgments on sentences containing long-distance dependencies were more sensitive to lexical content than linguists’ judgments were. Even sentences included in linguistic textbooks, which one might presume to have clear-cut judgments, in reality are judged gradiently by people Juzek (2024). Nonetheless, LMs’ language skills are often evaluated on binary grammaticality judgments on sentences Dentella et al. (2024, 2023); Warstadt et al. (2019).
The fact that human judgments are gradient can have profound consequences on evaluations. For instance, Dentella et al. (2023) compared humans and LMs against predetermined binary acceptability labels, reporting that LMs’ performance correlated poorly. However, comparing gradient perplexity measures with the same human judgments revealed a strong positive correlation Hu et al. (2024). Using perplexity measures for models (as well as allowing humans to provide ordinal or gradient judgments) is a step in the right direction Hu et al. (2024); Juzek (2024).
@@
Parsing text for universal dependencies (UD, de Marneffe et al., 2021) has become a well-established task for evaluating models Zeman et al. (2017, 2018), and since Hewitt and Manning (2019) showed BERT Devlin et al. (2019) to be somewhat skilled in UD, it has became the default operationalization of syntax in the NLP world Amini et al. (2023); Kryvosheieva and Levy (2025); Müller-Eberstein et al. (2022) and in discussions of inductive biases Lindemann et al. (2024); Glavaš and Vulić (2021). UD annotations are partially determined by semantics and they are based on lexical items, which makes them closer to the approach advocated here rather than abstract phrase structure rules. However, UD analyses presume a universal set of grammatical relations, which is problematic, because not all languages employ the same constructs. That is, there is no universally valid way to define or test for the syntax of nouns, verbs, adjectives, subjects, or direct objects (e.g., Croft, 2001). Moreover, UD requires an asymmetric relationship between a ’head’ and dependent, yet the long tail of language includes headless constructions (e.g., the Xer, the Yer construction) Michaelis (2003) and co-headed constructions (e.g., phrasal verbs, conjunctions, idioms). Therefore, UD annotations need to be determined for individual languages and need to allow for non-headed or co-headed cases to align well with formal aspects of natural languages.
@@ 3 Semantic Rules in LM Evaluation
Formal logic was developed as a branch of mathematics, used to prove mathematical and philosophical theorems and identify provability gaps Frege (1918); Russell (1905); Gödel (1931). It was based on algebraic rules operating on categorical and broadly defined categories. Notably, many logicians did not generally assume nor endorse using formal logic to represent the meanings of natural language utterances Carnap (1937); Baker and Hacker (1986), recognizing that natural languages differ from formal logic in many ways. [^1] Some logicians did advocate for using formal logic for natural language Tarski (1944); Montague (1970, 1973). For instance, logic treats and and but as equivalent. It does not provide a natural way to capture commands or questions (Austin, 1975), nor does it naturally distinguish presuppositions from assertions Strawson (1967). Finally, formal logic is not intended to capture effects of context Wittgenstein (1953); Russin et al. (2024).
Yet the assumption that natural language semantics can be modeled by formal logic has been made in the design of certain classic LM understanding benchmarks. Below, we review some instances and discuss their connection to formal semantics.
@@
Natural Language Inference tasks label the second of two sentences as an entailment, contradiction, or neutral, and this NLI task was originally used to train models Wang et al. (2019); Dagan et al. (2006); Nie et al. (2020). Today, NLI is used as a zero-shot evaluation metric to assess natural language understanding Zhou et al. (2024); McCoy et al. (2019). In introducing the Stanford NLI corpus, Bowman et al. (2015) state, “The semantic concepts of entailment and contradiction are central to all aspects of natural language meaning. ”) (emphasis added, see also Katz, 1972; van Benthem, 2008). While in the same paper, Bowman et al. (2015) acknowledge that judgments depend on many factors, such as commonsense knowledge, this fact is generally overlooked in papers that use NLI as a task to evaluate LMs’ general understanding.
Necessary and plausible inferences are a critical aspect of natural language understanding. However, they are highly dependent on the interlocutors’ general communicative goals. We aim to make sense of others’ messages, so we assume others are trying to be relevant and helpful and do our best to assign coherent meanings to all utterances Grice (1975). For example, outside of logic classes or heated arguments, people rarely conclude that two statements made by the same person are contradictory. If someone utters: ‘The boy is depressed but he is not DEPRESSED’, listeners do not throw up their hands and shout ‘contradiction!’. Instead, they may infer that the boy in question is only somewhat, and not extremely, depressed, or ask to learn more. People also assign interpretations to statements in ways that differ from what formal logic would predict (e.g., ‘run fast and you’ve got this’ or ‘If it snows, it snows.’ NLI tasks that rely on judging contradictions or entailments may over- or under-estimate how well LMs’ understanding of natural language aligns with humans’, particularly when binary judgments are required (cf. Dentella et al., 2024).
Evaluation metrics need to take humans’ communicative goals into account and allow for gradient and context-dependent interpretations. An example of the type of evaluation we endorse can be found in the underappreciated NOPE testbed. Parrish et al. (2021) selected 10 distinct constructions that trigger presuppositions and curated 100+ instances of each one, based on naturally occurring examples. Each stimulus includes two preceding sentences for context. The authors then collected gradient judgments from human raters, allowing them to use their ‘background knowledge about how the world works’ and compared the accuracy of several models, with several controls in place. This strikes us as a highly valuable blueprint for modern evaluations of LMs.
@@
Banarescu et al. (2013) introduced abstract graphical meaning representations (AMR) for sentential meaning that importantly includes aspects of lexical semantics. It was created to offer a repository of structured meanings to be used for evaluating understanding in LMs (Li et al., 2023; Qiu et al., 2022; Shaw et al., 2021, see also §4). Yet work on AMR concedes that it ignores so-called ‘syntactic idiosyncracies.’ For example, ‘he described her as a genius’ and ‘his description of her: genius’ are assigned the same AMR. Yet the former is unambiguously about her intellect, while the latter may instead be used to compliment his cleverly tactful description. More generally, simplifications of distinctions made in a natural language can be expected to result in lost meaning, since two utterances are rarely interchangeable in all contexts. Focusing on subtle but important differences in meaning offers an opportunity to design more challenging linguistic evaluations of LMs.